OpenStudy (anonymous):

Choose the correct simplification of the expression 3b/a^-2

OpenStudy (anonymous):

since I don't see much there to simplify, I'll only provide a hint.. \[\LARGE \frac{1}{x^y}=x^{-y}\] and vice versa... \[\LARGE \frac{1}{x^{-y}}=x^y\]

OpenStudy (phi):

They want you to get rid of \( a^{-2} \) in the denominator. The rule (as posted by Kreshnik) is "if you move the term to the top, multiply its exponent by -1" The exponent is the small number to the upper right